• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name adaq

    Log entry time 00:01:59 on April 21, 2010

    Entry number 318540

    keyword=Beam Time Accounting
    Hall A

    04/20/10 1600-2400
    Experiment #: E06-002
    Shift Leader: V. Sulkosky
    Shift Workers: M. Canan, J. Han Lee

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    0.000.000.000.000.920.630.970.723.23
    Beam Available
    but Not in Use
    0.000.000.000.000.000.140.000.250.39
    Beam Not Available
    or Unacceptable
    1.001.001.001.000.080.230.030.034.38
    Accelerator Con-
    figuration Change
    0.000.000.000.000.000.000.000.000.00
    Experiment
    Ready
    1.001.001.001.001.000.871.000.757.62
    Planned Config-
    uration Change
    0.000.000.000.000.000.130.000.000.13
    Unplanned
    Experiment Down
    0.000.000.000.000.000.000.000.250.25
    Hall not
    ready reasons
    Target0.000.000.000.000.000.000.000.000.00
    Data Acq0.000.000.000.000.000.000.000.250.25
    Other0.000.000.000.000.000.000.000.000.00
    HRSL0.000.000.000.000.000.000.000.000.00
    HRSR0.000.000.000.000.000.000.000.000.00
    HALL_OFF0.000.000.000.000.000.000.000.000.00

    Comments: 16:00 Waiting for beam restoration. 18:08-18:50 Tuning for Compton. 18:50-19:55 No beam due to BLM trips. 19:55-20:50 Compton Tuning.

    Data Entry by: V. Sulkosky